OpenStudy (tate_langdon):

Can someone check my work and show correct work Simplify the following expression. 2/3 ÷ 9/16 A. 3/8 B. 27/32 C. 1/6 D. 1 5/27

OpenStudy (tate_langdon):

I say the answer is D

OpenStudy (tate_langdon):

@sunnnystrong

jhonyy9 (jhonyy9):

how you get D. ? show your work please

OpenStudy (tate_langdon):

Tbh that's why i'm asking for the work weirdly my answers are right but my work is always wrong lol

jhonyy9 (jhonyy9):

you need getting the common denominator for 3 and 16

OpenStudy (sunnnystrong):

\[\frac{ 2 }{ 3 }\div \frac{ 9}{ 16 }=\frac{ 2 }{ 3 }*\frac{ 16 }{ 9 }=\frac{ 32 }{ 27 }=\] ... hmmmm

OpenStudy (tate_langdon):

D?

OpenStudy (sunnnystrong):

nah... is B written correctly?

OpenStudy (tate_langdon):

D has to be correct

jhonyy9 (jhonyy9):

sorry - there is divide not plus ?

OpenStudy (tate_langdon):

My teacher gave me back homework cause the answers were right but the work was wrong

OpenStudy (tate_langdon):

https://media.studyisland.com/cgi-bin/mimetex.cgi?1 \,%20\frac{5}{27}

OpenStudy (tate_langdon):

Maybe I wrote it wrong

OpenStudy (sunnnystrong):

no i don't think it would be b. you would be breaking math laws haha it's 32/27... but the option isn't there?

OpenStudy (sunnnystrong):

i mean d*

OpenStudy (tate_langdon):

diddnt I put D?

OpenStudy (sunnnystrong):

none of the options you listed are correct. it's e. 32/27 XD

OpenStudy (tate_langdon):

Lmao XD

OpenStudy (tate_langdon):

Yeah idk maybe my teacher had too much coffee?

OpenStudy (sunnnystrong):

happens to all of them.. if you just bring it up he will explain it to you.

OpenStudy (tate_langdon):

Yeah lol